OpenStudy (anonymous):

Simplify (sqrt) - 54 + (sqrt) - 96

OpenStudy (lalaly):

\[\sqrt{-54}+\sqrt{-96}?\]

OpenStudy (cwrw238):

\sqrt-54\

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

@lalaly Yes

OpenStudy (anonymous):

-54 = -9 x 6 root(54)= root (-9 x 6) = 3i x root6 similarly -96 = -16 x 6 root(96) = 4i x root 6

OpenStudy (anonymous):

Thats not correct.

OpenStudy (anonymous):

do I add them up?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7i (sqrt) 6?

OpenStudy (lalaly):

\[3i \sqrt{6}+4i \sqrt{6}=7i \sqrt{6}\]

OpenStudy (anonymous):

correct